OpenStudy (anonymous):

A sphere has a volume of 972π in.3. Find its radius.

OpenStudy (anonymous):

Hikari, you need to use the formula for the volume of a sphere:\[V=\frac{4}{3}\pi r^3\]You have the volume given to you already, so all you need to do is rearrange that formula to find the radius:\[r=\sqrt[3]{\frac{3V}{4 \pi}}=\sqrt[3]{\frac{3\times 972 \pi}{4 \pi}}=\sqrt[3]{729}=9\]So the radius is 9 inches.